OpenStudy (indiiavery):

Arrange these numbers from least to greatest. -3/7,-5/6,-3/5,-7/9 Answers: A.-5/6,-7/9,-3/7,-3/5 B.-3/7,-3/5,-7/9,-5/6 C.-3/7,-7/9,-3/5,-5/6 D.-7/9,-3/7,-5/6,-3/5

Mehek (mehek14):

lets change it to decimal form so it is easier

OpenStudy (anonymous):

what do you think:)?

OpenStudy (anonymous):

@IndiiAvery

Mehek (mehek14):

-3/7 = -0.42857142857 but lets round it to the nearest tenth -3/7 = -0.43 -5/6 = -0.83 -3/5 = -0.6 -7/9 = -0.78 remember the larger the negative number, the least it will be

OpenStudy (indiiavery):

Okay, and @ayanna1100 I think it's either B or C not quiet sure though.

OpenStudy (anonymous):

ok well @Mehek14 took the words right out of my mouth but b or c is a good answer :))))

Mehek (mehek14):

is -7/9 greater than -3/5?

OpenStudy (indiiavery):

So it would be answer B, right? @Mehek14

Mehek (mehek14):

yup :]

OpenStudy (indiiavery):

Thank you @Mehek14 & @ayanna1100 ! :D

OpenStudy (anonymous):

your wlcm always nice to help :))))